# Ratio Covers of Convex Sets and Optimal Mixture Density Estimation

### v1

@@ Appendix A High-probability online-to-batch conversion
We present a self-contained version of the argument of Theorem 1 in [53] specific to the logarithmic loss and a finite class of densities.
@@
The proof repeats the standard lines for the logarithmic loss [56]. See the proof of a more general Proposition 1 in [53]. ∎
@@
Let $T\subseteq\mathbb{R}^{2}_{+}$ and assume that for some $m>0$, $\left|\log\left(x/y\right)\right|\leqslant m$ for every $(x,y)\in T$. Then, for all $(x,y)\in T$,
@@
Let $h(w)=-\log w$. Then $h^{\prime}(w)=-w^{-1}$ and $h^{\prime\prime}(w)=w^{-2}=(h^{\prime}(w))^{2}$. Fix $(x,y)\in T$ and set $s=\frac{x+y}{2}$. Define
@@
Another useful lemma is a simple version of the Freedman’s inequality. For the version with these explicit constants we refer to [8, Theorem 1]. Throughout this section we denote ${\mathbf{E}}_{i-1}[\cdot]={\mathbf{E}}[\cdot|X_{1},\ldots,X_{i-1}]$.
@@
${\mathbf{E}}\log\left(\widehat{p}(X)\right)+{\mathbf{E}}\log\left(p_{j}(X)\right)\leqslant\frac{1}{n}\sum_{i=1}^{n}{\mathbf{E}}\left[{\mathbf{E}}_{i-1}\,r_{i}\right].$ (19)
@@
Inserting this into (19) and minimizing over $j\in[M]$, we have
@@
Appendix B Suboptimality of Bayesian model averaging
In this section, we show that our somewhat involved estimators might be required, as some of the most straightforward strategies fail to produce optimal error guarantees. In the convex aggregation setting, Theorem 2 of [44] shows that any estimator that does not depend on $\delta$ must incur a suboptimal high-probability error bound. In particular, there exist instances (already with $M=2$) for which any $\delta$-independent estimator, including the progressive mixture estimator of Yang and Barron (it is formally defined in Theorem˜7), incurs a lower bound of order
@@ Appendix B Suboptimality of Bayesian model averaging
which is worse by an extra $\log\log(1/\delta)$ factor than what is achieved by our $\delta$-dependent estimator in Theorem˜6.
In the remainder of this section, we show that an even stronger lower bound holds for Bayesian model averaging, which is arguably the most natural model aggregation procedure. For completeness, we also provide a matching upper bound for the case of two densities, showing that our lower-bound construction is tight up to absolute constants. Our lower bound focuses on two Bernoulli distributions (one of which is a point mass), as in the lower-bound construction of [44].
@@
$\operatorname{KL}(p^{\star},\widehat{p})\leqslant 5\,\operatorname{H}^{2}(p^{\star},\widehat{p})\,\sup_{x}\log\left(\frac{p^{\star}(x)}{\widehat{p}(x)}\right).$
Using (23) and the convexity of the squared Hellinger distance in its second argument, $\operatorname{H}^{2}(p^{\star},\widehat{p})\leqslant\operatorname{H}^{2}(p^{\star},q)$. We consider two cases. If $\operatorname{H}^{2}(p^{\star},q)\leqslant\frac{2\log(2/\delta)}{n}$, then
@@
Otherwise, $\operatorname{H}^{2}(p^{\star},q)>\frac{2\log(2/\delta)}{n}$ and we use $\operatorname{H}^{2}(p^{\star},\widehat{p})\leqslant 1$ to obtain
$\operatorname{KL}(p^{\star},\widehat{p})\leqslant 5\log\left(\frac{2}{\delta}\right).$
Combining the two bounds concludes the proof of the upper bound.
@@ Appendix C Birgé–Le Cam tournament
In this section we reproduce the details of the Birgé–Le Cam tournament procedure. The derivations are standard and we mainly follow the exposition in [47, Section 32.2] with the minor difference that we need to provide sharper tail bounds for our applications. For earlier results establishing this procedure we refer to [35, 12, 13]. For a set of densities $\mathcal{P}$, recall the definition (6) of the local Hellinger entropy. Let $\varepsilon_{n}>0$ satisfy the fixed-point inequality
@@
Theorem 9 (Le Cam-Birgé tournament).
@@
    $\mathcal{W}=\left\{q\in\mathcal{P}:\operatorname{H}(q,\widehat{p})\leqslant 21\varepsilon_{n}+8\sqrt{\frac{\log(2/\delta)}{n}}\right\}.$
@@
We first, formulate the following composite hypothesis testing result.
@@
$N_{\operatorname{H}}(B_{\operatorname{H}}(p,2^{k}\eta)\cap\mathcal{P},\eta/2)\leqslant N_{\mathsf{loc}}(\mathcal{P},\varepsilon)^{k+1}$
Proof of Theorem˜9.
@@ D.1 Proof of Lemma˜1
Let $h(x)=\frac{f(x)}{g(x)}$. By the standard representation we have
$\operatorname{KL}(f,g)=\int f\log\frac{f}{g}\,d\mu=\int g\bigl(h\log h-h+1\bigr)\,d\mu.$
@@
The inequality (26) can be shown by treating $h\leqslant e$ and $h>e$ separately and checking that $\frac{h\log h-h+1}{(\sqrt{h}-1)^{2}}$ is increasing on $(0,e]$ while $\frac{h\log h-h+1}{(\sqrt{h}-1)^{2}\log h}$ is decreasing on $e,\infty)$; both meet at value $(\sqrt{e}-1)^{-2}$ when $h=e$. Applying (26) with $h=f/g$ and $M=\max\left\{1,\ \sup_{x\in\mathcal{X}}\log\left(\frac{f(x)}{g(x)}\right)\right\}$, we obtain
$\operatorname{KL}(f,g)\leqslant\frac{M}{(\sqrt{e}-1)^{2}}\int g(\sqrt{h}-1)^{2}d\mu.$
Finally, observe that $\int g(\sqrt{h}-1)^{2}\,d\mu=\int(\sqrt{f}-\sqrt{g})^{2}d\mu=2\operatorname{H}^{2}(f,g)$. Substituting yields the claim. ∎
@@ D.2 Proof of Corollary˜2
To conclude the theorem, apply ˜2 with $K$ as given, $C=32$, and $B$ a $32$-ratio cover of $K$ of size at most $2^{8d}$ provided by Theorem˜1. This yields an $\varepsilon$-approximate Pareto curve of size
@@ D.4 Proof of Proposition˜3
First, we present the required estimator, which is simpler than the one we used in Theorem˜2.
@@
    $\mathcal{W}=\left\{\left(1-\frac{1}{n}\right)p+\frac{1}{nM}\sum_{q\in\mathcal{P}}q:\ p\in\mathcal{P}\right\}.$ (28)
  • Run a version of the progressive density estimator of Lemma˜2 over $\mathcal{W}$ to output the final density $\widehat{p}$.
@@
Observe that $\widetilde{p}\in\mathcal{W}$ and that for each $p,q\in\mathcal{W}$ we have
$\sup_{p,q\in\mathcal{W}}\sup_{x\in\mathcal{X}}\log\left(\frac{p(x)}{q(x)}\right)\leqslant\log\bigl(1+nM\bigr)\leqslant\log(2nM).$
@@
whenever $n\geqslant 2$. When $n=1$, we have $\mathcal{W}=\left\{\frac{1}{M}\sum\limits_{q\in\mathcal{P}}q\right\}$ and the claim follows since $\widehat{p}=\frac{1}{M}\sum\limits_{q\in\mathcal{P}}q$.∎
@@ D.5 Proof of Proposition˜4
We assume throughout that $n\geqslant 2$. When $n=1$, $\mathcal{W}$ contains a single element, namely $\mathcal{W}=\left\{\frac{1}{M}\sum_{q\in\mathcal{P}}q\right\},$ so $\widehat{p}$ is this deterministic density and the stated bound is immediate.
The proof exploits the curvature of the logarithmic loss on the ratio-bounded convex class $\mathcal{W}$ via an offset-process argument (see [39, 41, 55] for related derivations). Recall that $\widehat{p}$ is the maximum-likelihood estimator over $\mathcal{W}$, namely
$\widehat{p}\in\arg\min_{p\in\mathcal{W}}\widehat{L}_{n}(p),\qquad\widehat{L}_{n}(p)=-\frac{1}{n}\sum_{i=1}^{n}\log p(X_{i}),\qquad L(p)=-{\mathbf{E}}\log p(X),$
where $X\sim p^{\star}$ and $\mathbf{E}$} denotes expectation under $p^{\star}$. As in the proof of Proposition˜3, the smoothing in the definition of $\mathcal{W}$ implies that
$\sup_{p,q\in\mathcal{W}}\sup_{x\in\mathcal{X}}\Bigl|\log\Bigl(\frac{p(x)}{q(x)}\Bigr)\Bigr|\leqslant\log(2nM).$ (30)
@@
Let $p^{\star}_{\mathcal{W}}\in\arg\min_{p\in\mathcal{W}}L(p)$. For every $\lambda>0$ it holds that
${\mathbf{E}}\exp\Big(\lambda\big(L(\widehat{p})-L(p^{\star}_{\mathcal{W}})\big)\Big)\leqslant{\mathbf{E}}\exp\Big(\lambda\sup_{p\in\mathcal{W}}Z(p)\Big),$ (31)
@@
$Z(p)=\frac{1}{n}\sum_{i=1}^{n}\left[4\varepsilon_{i}\log\left(\frac{p^{\star}_{\mathcal{W}}(X_{i})}{p(X_{i})}\right)-\frac{1}{2(\log(2nM)+1)}\log^{2}\left(\frac{p^{\star}_{\mathcal{W}}(X_{i})}{p(X_{i})}\right)\right],$
@@
Define the log-density ratio $r_{p}(x)=\log\left(\frac{p^{\star}_{\mathcal{W}}(x)}{p(x)}\right).$ Let $(X_{1}^{\prime},\dots,X_{n}^{\prime})$ be an independent ghost sample (i.i.d. as $(X_{i})$) and set
@@
By convexity of $\mathcal{W}$, the midpoint $\bar{p}=\frac{\widehat{p}+p^{\star}_{\mathcal{W}}}{2}$ belongs to $\mathcal{W}$. By the definition of $\widehat{p}$, it holds that
$\widehat{L}_{n}(\widehat{p})\leqslant\widehat{L}_{n}(\bar{p}).$ (32)
@@
$\log(\bar{p}(X_{i}))=-\log\Big(\tfrac{\widehat{p}(X_{i})+p^{\star}_{\mathcal{W}}(X_{i})}{2}\Big)\leqslant\frac{-\log\widehat{p}(X_{i})-\log p^{\star}_{\mathcal{W}}(X_{i})}{2}-\frac{1}{8(\log(2nM)+1)}\,r_{\widehat{p}}(X_{i})^{2}.$
@@
We repeat the same lines at the population level. Fix any $p\in\mathcal{W}$ and let $\widetilde{p}=(p+p^{\star}_{\mathcal{W}})/2\in\mathcal{W}$. Because $p^{\star}_{\mathcal{W}}$ is a population minimizer over $\mathcal{W}$, we have
@@
Applying Lemma 7 as above, using (34), taking the expectation and rearranging we obtain that for all $p\in\mathcal{W}$,
@@
We have $L(\widehat{p})-L(p^{\star}_{\mathcal{W}})={\mathbf{E}}r_{\widehat{p}}(X)$. For brevity of notation, set $\alpha=\frac{1}{4(\log(2nM)+1)}.$ The following lines of inequalities hold (conditioned on $\widehat{p}$):
@@
$\leqslant\sup_{p\in\mathcal{W}}\Big\{2\big({\mathbf{E}}-P_{n}\big)r_{p}-2\alpha\,P_{n}\big(r_{p}^{2}\big)-\alpha\,{\mathbf{E}}\big[r_{p}(X)^{2}\big]\Big\}$
$\leqslant\sup_{p\in\mathcal{W}}\Big\{2\big({\mathbf{E}}-P_{n}\big)r_{p}-\alpha P_{n}\big(r_{p}^{2}\big)-\alpha{\mathbf{E}}\big[r_{p}(X)^{2}\big]\Big\}.$
@@
${\mathbf{E}}\exp\Big(\lambda\big(L(\widehat{p})-L(p^{\star}_{\mathcal{W}})\big)\Big)$ $\leqslant{\mathbf{E}}\exp\big(\lambda\sup_{p\in\mathcal{W}}\Big\{2\big({\mathbf{E}}-P_{n}\big)r_{p}-\alpha\,P_{n}\big(r_{p}^{2}\big)-\alpha\,{\mathbf{E}}\big[r_{p}(X)^{2}\big]\Big\}\big)$
$\leqslant{\mathbf{E}}\exp\big(\lambda\sup_{p\in\mathcal{W}}\Big\{2\big(P_{n}^{\prime}-P_{n}\big)r_{p}-{\alpha}P_{n}\big(r_{p}^{2}\big)-{\alpha}P_{n}^{\prime}\big(r_{p}^{2}\big)\Big\}\big),$
$={\mathbf{E}}\exp\Big(\lambda\sup_{p\in\mathcal{W}}\frac{1}{n}\sum_{i=1}^{n}\Big[2\varepsilon_{i}\big(r_{p}(X_{i}^{\prime})-r_{p}(X_{i})\big)-\alpha\big(r_{p}(X_{i})^{2}+r_{p}(X_{i}^{\prime})^{2}\big)\Big]\Big)$
$\leqslant{\mathbf{E}}\exp\Big(\lambda\sup_{p\in\mathcal{W}}\frac{1}{n}\sum_{i=1}^{n}\Big[4\varepsilon_{i}r_{p}(X_{i})-2\alpha r_{p}(X_{i})^{2}\Big]\Big),$
@@
The next lemma is about the Lipschitzness of the loss class indexed by $\mathcal{W}$. For $\alpha\in\Delta_{M-1}$, we denote
@@
so that $\mathcal{W}=\{w_{\alpha}:\alpha\in\Delta_{M-1}\}$, where $\Delta_{M-1}$ is the simplex in $\mathbb{R}^{M}$.

### v2

@@ Appendix A High-probability online-to-batch conversion
We present a self-contained version of the argument of Theorem 1 in [56], adapted to the logarithmic loss and a finite class of densities.
@@
The proof repeats the standard lines for the logarithmic loss [59]. See the proof of the more general Proposition 1 in [56]. ∎
@@
Let $m>0$ and $x,y\in\mathbb{R}^{+}$ such that $\left|\log\left(x/y\right)\right|\leqslant m$. Then,
@@
Let $h(w)=-\log w$. Then $h^{\prime}(w)=-w^{-1}$ and $h^{\prime\prime}(w)=w^{-2}=(h^{\prime}(w))^{2}$. Fix $(x,y)\in\mathbb{R}^{2}$ and set $s=\frac{x+y}{2}$. Define
@@
Another useful lemma is a simple version of Freedman’s inequality. For the version with these explicit constants we refer to [10, Theorem 1]. Throughout this section we denote ${\mathbf{E}}_{i-1}[\cdot]={\mathbf{E}}[\cdot|X_{1},\ldots,X_{i-1}]$.
@@
${\mathbf{E}}\log\left(\widehat{p}(X)\right)+{\mathbf{E}}\log\left(p_{j}(X)\right)\leqslant\frac{1}{n}\sum_{i=1}^{n}{\mathbf{E}}_{i-1}\left[r_{i}\right].$ (23)
@@
Inserting this into (23) and finally choosing $j\in[M]$ that minimizes $\left(-{\mathbf{E}}\log\left(p_{j}(X)\right)\right)$, we have
@@
Appendix B Suboptimality of MLE and Bayesian model averaging
In this section, we show that simple baseline methods for model aggregation, such as MLE and Bayes model averaging, fail to achieve optimal guarantees in the distribution-free setting.
The following simple fact, mentioned in the introduction, shows that in the distribution-free setting, no selection procedure (such as MLE) achieves any nontrivial guarantee.
Fact 1.
For every $n\geqslant 1$, there exists a dictionary $\mathcal{P}=\{p_{1},p_{2}\}$ such that the following holds. For every estimator $\widehat{p}=\widehat{p}_{n}$ based on an i.i.d. sample $X_{1},\dots,X_{n}$ of size $n$ and taking values inside $\mathcal{P}$, there exists $p^{\star}\in\mathcal{P}$ such that if $X_{1},\dots,X_{n}\sim p^{\star}$, then $\mathbf{P}(\operatorname{KL}({p^{\star}},{\widehat{p}_{n}})=+\infty)\geqslant 0.99$.
Proof.
Let $\mathcal{X}=\{0,1,2\}$ and $\eta>0$ be such that $(1-\eta)^{n}=0.99$. Define the densities $p_{1},p_{2}$ on $\mathcal{X}$ (with respect to the counting measure) by $p_{1}(0)=1-\eta,p_{1}(1)=\eta,p_{1}(2)=0$ and $p_{2}(0)=1-\eta,p_{2}(1)=0,p_{2}(2)=\eta$. Let $E$ denote the event $\{X_{1}=\dots=X_{n}=0\}$ and $q$ the value taken by the estimator $\widehat{p}_{n}$ under $E$, such that (for both $p^{\star}=p_{1}$ and $p^{\star}=p_{2}$) $\mathbf{P}(\widehat{p}_{n}=q)\geqslant\mathbf{P}(E)=(1-\eta)^{n}=0.99$. Now if $q=p_{1}$ we take $p^{\star}=p_{2}$, and if $q=p_{2}$ we take $p^{\star}=p_{1}$; in both cases, $\operatorname{KL}({p^{\star}},{q})=+\infty$ and thus $\mathbf{P}(\operatorname{KL}({p^{\star}},{\widehat{p}_{n}})=+\infty)\geqslant 0.99$. ∎
Next, in the convex aggregation setting, Theorem 2 of [47] shows that any estimator that does not depend on $\delta$ must incur a suboptimal high-probability error bound. In particular, there exist instances (already with $M=2$) for which any expectation-optimal estimator, including the progressive mixture estimator of Yang and Barron (formally defined in Theorem˜7), incurs a lower bound of order
@@ Appendix B Suboptimality of MLE and Bayesian model averaging
which is worse than what is achieved by our $\delta$-dependent estimator in Theorem˜6.
In the remainder of this section, we show that an even stronger lower bound holds for Bayesian model averaging, which is arguably the most natural model aggregation procedure. For completeness, we also provide a matching upper bound for the case of two densities, showing that our lower bound is tight up to absolute constants. Our lower bound focuses on two Bernoulli distributions, as in the lower-bound construction of [47].
@@
$\operatorname{KL}(p^{\star},\widehat{p})\leqslant 5\operatorname{H}^{2}(p^{\star},\widehat{p})\log\left(\frac{2}{\delta}\right).$ (28)
We consider the cases. First, if $2\log(2/\delta)/n>1$, then the minimum in the statement is achieved at the second term and since $\operatorname{H}^{2}(p^{\star},\widehat{p})\leqslant 1$, we have
$\operatorname{KL}(p^{\star},\widehat{p})\leqslant 5\log\left(\frac{2}{\delta}\right),$
which proves the claim. Otherwise, we have $2\log(2/\delta)/n\leqslant 1$, and we again consider two cases. In the first case, assume $\operatorname{H}^{2}(p^{\star},q)\leqslant\frac{2\log(2/\delta)}{n}$. Using (27) and the convexity of the squared Hellinger distance in its second argument, $\operatorname{H}^{2}(p^{\star},\widehat{p})\leqslant\operatorname{H}^{2}(p^{\star},q)$, which implies
@@
Otherwise, we have $\operatorname{H}^{2}(p^{\star},q)>\frac{2\log(2/\delta)}{n}$. In this case we want to upper bound $\widehat{\alpha}$. We have as in the proof of Lemma˜3 for any $t>0$,
$\mathbf{P}\left(\frac{\prod_{i=1}^{n}q(X_{i})}{\prod_{i=1}^{n}p^{\star}(X_{i})}\geqslant t\right)\leqslant\frac{1}{\sqrt{t}}\bigl(1-\operatorname{H}^{2}(p^{\star},q)\bigr)^{n}\leqslant\frac{1}{\sqrt{t}}\exp\bigl(-n\operatorname{H}^{2}(p^{\star},q)\bigr).$
We choose $t=\frac{2\log(2/\delta)}{n\operatorname{H}^{2}(p^{\star},q)}$. Note that $t<1<1/\delta$. By the above line we have
$\mathbf{P}\left(\frac{\prod_{i=1}^{n}q(X_{i})}{\prod_{i=1}^{n}p^{\star}(X_{i})}\geqslant\frac{2\log(2/\delta)}{n\operatorname{H}^{2}(p^{\star},q)}\right)\leqslant\sqrt{\frac{n\operatorname{H}^{2}(p^{\star},q)}{2\log(2/\delta)}}\exp(-n\operatorname{H}^{2}(p^{\star},q))\leqslant\exp(-2\log(2/\delta))\leqslant\delta,$
where we used that $\sqrt{x}\exp(-x)$ is decreasing for $x\geqslant 1/2$ and $n\operatorname{H}^{2}(p^{\star},q)>2\log(2/\delta)>1/2$. Therefore, on the complementary event, $\widehat{\alpha}\leqslant\frac{\prod_{i=1}^{n}q(X_{i})}{\prod_{i=1}^{n}p^{\star}(X_{i})}<\frac{2\log(2/\delta)}{n\operatorname{H}^{2}(p^{\star},q)},$ and (27) also holds since $t<1/\delta$. Hence, by (28) and the convexity of the squared Hellinger distance in its second argument, we have on this event
$\operatorname{KL}(p^{\star},\widehat{p})\leqslant 5\operatorname{H}^{2}(p^{\star},\widehat{p})\log\left(\frac{2}{\delta}\right)\leqslant 5\widehat{\alpha}\operatorname{H}^{2}(p^{\star},q)\log\left(\frac{2}{\delta}\right)\leqslant\frac{10\log^{2}(2/\delta)}{n}.$
The upper bound follows.
@@ Appendix C Birgé–Le Cam tournament
In this section we reproduce the details of the Birgé–Le Cam tournament procedure. The derivations are standard and we mainly follow the exposition in [50, Section 32.2] with the minor difference that we need to provide sharper tail bounds for our applications. For earlier results establishing this procedure we refer to [38, 14, 15]. For a set of densities $\mathcal{P}$, recall the definition (10) of the local Hellinger entropy. Let $\varepsilon_{n}>0$ satisfy the fixed-point inequality
@@
Lemma 9 (Le Cam-Birgé tournament).
@@
    $\widehat{\mathcal{Q}}_{1}=\left\{q\in\mathcal{P}:\operatorname{H}(q,\widehat{p})\leqslant 21\varepsilon_{n}+8\sqrt{\frac{\log(2/\delta)}{n}}\right\}.$
@@
We first recall the following result on hypotheses testing between Hellinger balls, due to Le Cam [38] and Birgé [14].
@@
$N_{\operatorname{H}}(B_{\operatorname{H}}(p,2^{k}\eta)\cap\mathcal{P},\eta/2)\leqslant N_{\mathsf{loc}}(\mathcal{P},\varepsilon)^{k+1}.$
Proof of Lemma˜9.
@@ D.1 Proof of Lemma˜1
Let h(x)=p(x)/q(x)h(x)={p(x)}/{q(x)}. By the standard representation we have
$\operatorname{KL}(p,q)=\int p\log\frac{p}{q}\,d\mu=\int q\bigl(h\log h-h+1\bigr)\,d\mu.$
@@
The inequality (31) can be shown by treating $h\leqslant e$ and $h>e$ separately and checking that $\frac{h\log h-h+1}{(\sqrt{h}-1)^{2}}$ is increasing on $(0,e]$ while $\frac{h\log h-h+1}{(\sqrt{h}-1)^{2}\log h}$ is decreasing on $e,\infty)$; both meet at value $(\sqrt{e}-1)^{-2}$ when $h=e$. Applying (31) with $h=p/q$ and $m=\max\left\{1,\ \sup_{x\in\mathcal{X}}\log\left(\frac{p(x)}{q(x)}\right)\right\}$, we obtain
$\operatorname{KL}(p,q)\leqslant\frac{m}{(\sqrt{e}-1)^{2}}\int q(\sqrt{h}-1)^{2}d\mu.$
Finally, observe that $\int q(\sqrt{h}-1)^{2}\,d\mu=\int(\sqrt{p}-\sqrt{q})^{2}d\mu=2\operatorname{H}^{2}(p,q)$. Substituting yields the claim. ∎
@@ D.2 Proof of Corollary˜2
To conclude the theorem, apply ˜2 with $K$ as given, $C=32$, and $B$ a $32$-ratio cover of $K$ of size at most $2^{8d}$ provided by Theorem˜3. This yields an $\varepsilon$-approximate Pareto curve of size
@@ D.4 Proof of Proposition˜3
First, we present the required estimator, which is simpler than the one we used in Theorem˜1.
@@
    $\mathcal{Q}=\left\{\left(1-\frac{1}{n}\right)p+\frac{1}{nM}\sum_{q\in\mathcal{P}}q:\ p\in\mathcal{P}\right\}.$ (33)
  • Run the density estimator of Lemma˜2 over $\mathcal{Q}$ to output the final density $\widehat{p}$.
@@
Observe that $\widetilde{p}\in\mathcal{Q}$ and that for each $p,q\in\mathcal{Q}$ we have
$\sup_{p,q\in\mathcal{Q}}\sup_{x\in\mathcal{X}}\log\left(\frac{p(x)}{q(x)}\right)\leqslant\log\bigl(1+nM\bigr)\leqslant\log(2nM).$
@@
whenever $n\geqslant 2$. When $n=1$, we have $\mathcal{Q}=\big\{\frac{1}{M}\sum\limits_{q\in\mathcal{P}}q\big\}$ and the claim follows since $\widehat{p}=\frac{1}{M}\sum\limits_{q\in\mathcal{P}}q$.∎
@@ D.5 Proof of Proposition˜4
We assume throughout that $n\geqslant 2$. When $n=1$, $\mathcal{Q}$ contains a single element, namely $\mathcal{Q}=\left\{\frac{1}{M}\sum_{q\in\mathcal{P}}q\right\},$ so $\widehat{p}$ is this deterministic density and the stated bound is immediate.
The proof exploits the curvature of the logarithmic loss on the ratio-bounded convex class $\mathcal{Q}$ via an offset-process argument (see [42, 44, 58] for related derivations). Recall that $\widehat{p}$ is the maximum-likelihood estimator over $\mathcal{Q}$, namely
$\widehat{p}\in\arg\min_{p\in\mathcal{Q}}\widehat{L}(p),\qquad\widehat{L}(p)=-\frac{1}{n}\sum_{i=1}^{n}\log p(X_{i}),\qquad L(p)=-{\mathbf{E}}\log p(X),$
where $X\sim p^{\star}$ and $\mathbf{E}$} denotes expectation under $p^{\star}$. As in the proof of Proposition˜3, the smoothing in the definition of $\mathcal{Q}$ implies that
$\sup_{p,q\in\mathcal{Q}}\sup_{x\in\mathcal{X}}\Bigl|\log\Bigl(\frac{p(x)}{q(x)}\Bigr)\Bigr|\leqslant\log(2nM).$ (35)
@@
Let $p^{\star}_{\mathcal{Q}}\in\arg\min_{p\in\mathcal{Q}}L(p)$. For every $\lambda>0$ it holds that
${\mathbf{E}}\exp\Big(\lambda\big(L(\widehat{p})-L(p^{\star}_{\mathcal{Q}})\big)\Big)\leqslant{\mathbf{E}}\exp\Big(\lambda\sup_{p\in\mathcal{Q}}Z(p)\Big),$ (36)
@@
$Z(p)=\frac{1}{n}\sum_{i=1}^{n}\left[4\varepsilon_{i}\log\left(\frac{p^{\star}_{\mathcal{Q}}(X_{i})}{p(X_{i})}\right)-\frac{1}{2(\log(2nM)+1)}\log^{2}\left(\frac{p^{\star}_{\mathcal{Q}}(X_{i})}{p(X_{i})}\right)\right],$
@@
Define the log-density ratio $r_{p}(x)=\log\left(\frac{p^{\star}_{\mathcal{Q}}(x)}{p(x)}\right).$ Let $(X_{1}^{\prime},\dots,X_{n}^{\prime})$ be an independent ghost sample (i.i.d. as $(X_{i})$) and set
@@
By convexity of $\mathcal{Q}$, the midpoint $\bar{p}=\frac{\widehat{p}+p^{\star}_{\mathcal{Q}}}{2}$ belongs to $\mathcal{Q}$. By the definition of $\widehat{p}$, it holds that
$\widehat{L}(\widehat{p})\leqslant\widehat{L}(\bar{p}).$ (37)
@@
$\log(\bar{p}(X_{i}))=-\log\Big(\tfrac{\widehat{p}(X_{i})+p^{\star}_{\mathcal{Q}}(X_{i})}{2}\Big)\leqslant\frac{-\log\widehat{p}(X_{i})-\log p^{\star}_{\mathcal{Q}}(X_{i})}{2}-\frac{1}{8(\log(2nM)+1)}\,r_{\widehat{p}}(X_{i})^{2}.$
@@
We repeat the same lines at the population level. Fix any $p\in\mathcal{Q}$ and let $\widetilde{p}=(p+p^{\star}_{\mathcal{Q}})/2\in\mathcal{Q}$. Because $p^{\star}_{\mathcal{Q}}$ is a population minimizer over $\mathcal{Q}$, we have
@@
Applying Lemma 7 as above, using (39), taking the expectation and rearranging we obtain that for all $p\in\mathcal{Q}$,
@@
We have $L(\widehat{p})-L(p^{\star}_{\mathcal{Q}})={\mathbf{E}}r_{\widehat{p}}(X)$. For brevity of notation, set $\alpha=\frac{1}{4(\log(2nM)+1)}.$ The following lines of inequalities hold (conditioned on $\widehat{p}$):
@@
$\leqslant\sup_{p\in\mathcal{Q}}\Big\{2\big({\mathbf{E}}-P_{n}\big)r_{p}-2\alpha\,P_{n}\big(r_{p}^{2}\big)-\alpha\,{\mathbf{E}}\big[r_{p}(X)^{2}\big]\Big\}$
$\leqslant\sup_{p\in\mathcal{Q}}\Big\{2\big({\mathbf{E}}-P_{n}\big)r_{p}-\alpha P_{n}\big(r_{p}^{2}\big)-\alpha{\mathbf{E}}\big[r_{p}(X)^{2}\big]\Big\}.$
@@
${\mathbf{E}}\exp\Big(\lambda\big(L(\widehat{p})-L(p^{\star}_{\mathcal{Q}})\big)\Big)$ $\leqslant{\mathbf{E}}\exp\big(\lambda\sup_{p\in\mathcal{Q}}\Big\{2\big({\mathbf{E}}-P_{n}\big)r_{p}-\alpha\,P_{n}\big(r_{p}^{2}\big)-\alpha\,{\mathbf{E}}\big[r_{p}(X)^{2}\big]\Big\}\big)$
$\leqslant{\mathbf{E}}\exp\big(\lambda\sup_{p\in\mathcal{Q}}\Big\{2\big(P_{n}^{\prime}-P_{n}\big)r_{p}-{\alpha}P_{n}\big(r_{p}^{2}\big)-{\alpha}P_{n}^{\prime}\big(r_{p}^{2}\big)\Big\}\big),$
$={\mathbf{E}}\exp\Big(\lambda\sup_{p\in\mathcal{Q}}\frac{1}{n}\sum_{i=1}^{n}\Big[2\varepsilon_{i}\big(r_{p}(X_{i}^{\prime})-r_{p}(X_{i})\big)-\alpha\big(r_{p}(X_{i})^{2}+r_{p}(X_{i}^{\prime})^{2}\big)\Big]\Big)$
$\leqslant{\mathbf{E}}\exp\Big(\lambda\sup_{p\in\mathcal{Q}}\frac{1}{n}\sum_{i=1}^{n}\Big[4\varepsilon_{i}r_{p}(X_{i})-2\alpha r_{p}(X_{i})^{2}\Big]\Big),$
@@
The next lemma is about the Lipschitzness of the loss class indexed by $\mathcal{Q}$. For $\alpha\in\Delta_{M-1}$, we denote
@@
so that $\mathcal{Q}=\{w_{\alpha}:\alpha\in\Delta_{M-1}\}$, where $\Delta_{M-1}$ is the simplex in $\mathbb{R}^{M}$.